OpenStudy (anonymous):

One more question about matrices and I'm done!! Sorry! ._. I need to figure out how to solve matrix equations. I'm confuzzled.

OpenStudy (anonymous):

For example, I need to solve this one:\[\left[\begin{matrix}4 & 8 \\ 2 & 5\end{matrix}\right]\left(\begin{matrix}x \\ y\end{matrix}\right)=\left(\begin{matrix}0 \\ 6\end{matrix}\right)\]

OpenStudy (dan815):

do u know how matrix multiplication works , or a dot product

OpenStudy (dan815):

another way is to take inverse of both sides, prolly not for ur lvl yet

OpenStudy (anonymous):

This is all really new to me /: If anyone could walk me through just this one problem step by step it would probably give me a better understanding.

OpenStudy (dan815):

4 * x + 8* y = 0 2*x + 5*y = 6 are ur 2 equations to solve

OpenStudy (anonymous):

Is it possible to solve those by putting them into a graphing calculator and finding the intersecting point?

OpenStudy (dan815):

[4 8 |0] [2 5 |6] get this in this form below [1 0 | x] [0 1 | y]

OpenStudy (dan815):

or just solve those 2 equations above from basic algebra u have 2 functions and 2 unknown

OpenStudy (dan815):

4x + 8y = 0 ; x =-8y/4 =-2y ; sub x in other equation and solve 2x + 5y = 6

OpenStudy (dan815):